# CVBench: Benchmarking Cross-Video Synergies for Complex Multimodal Reasoning

### v3

@@ Abstract
Abstract
While multimodal large language models (MLLMs) exhibit strong performance on single-video tasks (e.g., video question answering), their ability across multiple videos remains critically underexplored. However, this capability is essential for real-world applications, including multi-camera surveillance and cross-video procedural learning. To bridge this gap, we present CVBench, the first comprehensive benchmark designed to assess cross-video relational reasoning rigorously. CVBench comprises 1,000 question-answer pairs spanning three hierarchical tiers: cross-video object association (identifying shared entities), cross-video event association (linking temporal or causal event chains), and cross-video complex reasoning (integrating commonsense and domain knowledge). Built from five domain-diverse video clusters (e.g., sports, life records), the benchmark challenges models to synthesise information across dynamic visual contexts. Extensive evaluation of 10+ leading MLLMs (including GPT-4o, Gemini-2.0-flash, Qwen2.5-VL) under zero-shot or chain-of-thought prompting paradigms. Key findings reveal stark performance gaps: even top models, such as GPT-4o, achieve only 60% accuracy on causal reasoning tasks, compared to the 91% accuracy of human performance. Crucially, our analysis reveals fundamental bottlenecks inherent in current MLLM architectures, notably deficient inter-video context retention and poor disambiguation of overlapping entities. CVBench establishes a rigorous framework for diagnosing and advancing multi-video reasoning, offering architectural insights for next-generation MLLMs.The data and evaluation code are available at https://github.com/Hokhim2/CVBench.
keywords:
Multimodal large language models, Cross-video reasoning, Video question answering, Multivideo understanding, Benchmark evaluation.
journal: Pattern Recognition
[…]
[label1]organization=School of Systems Science and Engineering, Sun Yat-sen University,addressline=No. 135 Xingang West Road, city=Guangzhou, postcode=510275, state=Guangdong, country=China
[…]
[label2]organization=Tencent,addressline=Keji Zhongyi Road, Nanshan District, city=Shenzhen, postcode=518057, state=Guangdong, country=China
[…]
[label3]organization=School of Electronics and Communication Engineering, Sun Yat-sen University,addressline=No. 66 Gongchang Road, city=Shenzhen, postcode=518107, state=Guangdong, country=China
[…]
[label4]organization=School of Electronics and Information Engineering, Foshan University,addressline=No. 18 Jiangwan 1st Road, city=Foshan, postcode=528225, state=Guangdong, country=China
[…]
[label5]organization=Department of Computer Science and Engineering, Southern University of Science and Technology, addressline=1088 Xueyuan Avenue, city=Shenzhen, postcode=518055, state=Guangdong, country=China
{highlights}
First systematic benchmark for evaluating spatiotemporal relational pattern recognition across multiple videos, featuring a hierarchical structure with 1,000 expert-annotated QA pairs to diagnose capabilities in object association, event association, and complex reasoning.
Comprehensive evaluation reveals significant limitations of current models: even state-of-the-art multimodal models achieve only 63.5% accuracy on causal reasoning, far below human performance of 91.3%, exposing fundamental bottlenecks in inter-video pattern integration and entity disambiguation.
Provides actionable architectural insights for advancing pattern recognition models, including cross-video memory mechanisms and grounded spatiotemporal reasoning, offering a clear pathway toward robust multi-video understanding systems.
@@ 1 Introduction
The ability to reason across multiple video streams is fundamental to real-world applications, such as multi-camera surveillance systems, distributed procedural learning, and multi-view activity analysis Wu2025 . While multimodal large language models (MLLMs) have demonstrated significant progress in single-video understanding CAI2025111670 , their capacity for spatiotemporal pattern reasoning across videos remains a critical and underexplored challenge in pattern recognition research ZHENG2025111319 . This capability requires models to identify and analyze relational patterns across distinct spatiotemporal contexts 11008449 , such as entity correspondences, event sequences, and causal links. The research gap stems not merely from a lack of single-video benchmarks, but from the inherent limitations of existing evaluations in capturing cross-video relational dynamics, which are central to real-world pattern recognition systems.
Current MLLMs face fundamental barriers in cross-video scenarios: 1) Inter-video context linking: models lack mechanisms to persistently integrate and retain object/event states cross-video separated by time and space Li10658165 , 11093290 ; 2) Entity disambiguation: overlapping or visually similar entities cross-video lack unique spatiotemporal grounding, causing identification errors; 3) Temporal-causal modeling: existing architectures poorly capture long-range dependencies and causal links spanning asynchronous video sequences 11095110 .
Figure 1: Illustration of CVBench task definition. CVBench focuses on three core tasks that are crucial for real-world applications: a) cross-video object association (identifying shared entities across different videos), b) cross-video event association (linking temporal or causal event chains that span across video sequences), and c) cross-video complex reasoning (integrating commonsense to understand and reason about complex interrelationships between video content). The benchmark format is multi-choice QA, and here we only show the correct answer for visualization.
@@
1) We introduce CVBench, the first diagnostic benchmark explicitly targeting cross-video understanding through object association, event association, and complex reasoning.
2) We release a carefully curated multi-domain dataset with comprehensive annotations, enabling rigorous stress-testing of spatiotemporal integration and multi-view synthesis capabilities.
3) We provide a systematic analysis of state-of-the-art MLLMs on CVBench, exposing critical performance gaps and identifying key research directions for advancing cross-video pattern reasoning in the field.
@@ 2 Related work
Multimodal large language models
The evolution of MLLMs has progressed from image-based Zhang10547418 to video-capable architectures Fu11093290 , ZHANG2026111925 , marked by two main improvements: 1) Long-range context modeling that enables efficient handling excessively long temporal contexts Yue11093104 , Song10657734 ; 2) Spatio-temporal fusion that integrates both appearance and motion dynamics with language reasoning Ren10656135 , Zhang11094660 , Liu11094049 . Despite these advances, current MLLMs remain architecturally constrained to single-video scenario, lacking mechanisms for cross-video interactions, exhibiting inadequate spatiotemporal grounding for overlapping entities, and demonstrating limited modeling of temporal event chains. These fundamental limitations restrict their applicability to multi-camera surveillance and cross-video applications Yang11037495 .
Video understanding benchmarks
A comprehensive and objective benchmark is essential for evaluating MLLMs, enabling the comparison and investigation of the performance of various models SHAO2026112724 . Recent advances in multimodal learning have led to the development of several benchmarks for video understanding tasks ZHONG2025111035 . However, existing video evaluation frameworks exhibit critical limitations in assessing cross-video reasoning capabilities. First, general multimodal frameworks are predominantly with broad domain coverage, such as Video-MMMU and Video-MME Fu11093290 ), to evaluate instruction-following capabilities but lacking cross-video relational tasks. Second, long-term video benchmarks contain longer videos with multiple sub-events, however, are confined to single source or similar creators, such as VideoTree Wang11094563 , Video-Bench Ning11284911 , and Chapter-Llama Ventura11092338 . These deficiencies fundamentally limit their applicability to real-world scenarios in multi-camera surveillance and cross-video procedural learning.
@@ 3.1 Task definition
1) Object association: Cross-video object recognition, multi-video attribute recognition, joint-video counting, and cross-video entity matching. These require consistent identification of entities across domains or viewpoints despite occlusion or deformation.
2) Event association: Cross-video anomaly detection, scene recognition, key-action recognition, event retrieval, multi-view scene understanding, and temporal reasoning. These demand fusion of spatiotemporal signatures across videos to reconstruct event dynamics.
3) Complex reasoning: Joint-video spatial navigating, video difference captioning, cross-video counterfactual reasoning, joint-video summarization, and procedural transfer. These necessitate higher-order synthesis including spatial imagination, fine-grained reasoning, causal inference and adaptive knowledge transfer.
@@ 3.2 Dataset collection
3.2.1 Video curation
We employ task-driven sampling from YouTube to construct diverse video groups (at most four videos per group). Detailed statistics are shown in Fig. 2. We ensure source diversity through the inclusion of multi-creator channels (videos without similar or same creators), cross-domain associations (videos across different domain but with shared topics or intrinsic relations), and the segmentation of long videos (sequential occurrence of event-level activities). Specifically, we measure video quality under the following four curation criteria: a) Multi-domain coverage, including daily life, sports, surveillance and other commen scenrios; b) Intrinsic inter-video relationships, such as differences, similarities, co-occurrence, causality causal, temporal order; c) Duration constraints, where videos are primarily within 10 minutes in length, and those exceeding 10 minutes account for less than 5% of the dataset; d) Visual quality and availability, requiring videos to have a resolution of at least 720p and a frame rate of 24 fps.
@@ 3.3 Comparison with previous benchmarks
Benchmark #Video #QA #Video Per QA Duration (s) Multi-level CV.U CV.R
TempCompass tempcompass2024 410 500 1 11.4
MSVD-QA XuVideo 504 13,157 1 9.8
Video-MME Fu11093290 900 2,700 1 1017.9
NExT-QA Xiao9577425 1,000 8,564 1 39.5
AutoEval-Video chen2024Auto 327 327 1 14.6
EgoSchema egoschema2023 5,063 5,063 1 180.0
ActivityNet-QA zhou2019Activity 800 8,000 1 111.4
CVBench (Ours) 1,315 1,000 2 $\sim$ 4 106.6
4 Benchmarked results
4.1 Experimental settings
4.1.1 Models
Our evaluation comprehensively assesses various MLLMs, including closed-source and open-source models. A prominent model in our suite is Gemini team2023gemini , a closed-source model known for its sophisticated multimodal processing capabilities. On the open-source side, Qwen2.5-VL videor1 stands out, offering transparency with its accessible codebase for further exploration and customization. We also evaluate advanced multimodal reasoning models such as Video-R1 videor1 , which leverages reinforcement learning to integrate multimodal information and perform complex cross-video reasoning tasks. For ablation study, we choose Qwen2.5VL-7B as the default model.
4.1.2 Settings
@@
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport 52.4 70.9 22.7 31.0 56.4 57.7 75.0 60.8
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport 51.3 80.0 22.7 26.2 50.9 55.8 69.2 60.8
LLaVA-OneVision-7B Li2024LLamavid 52.6 83.6 40.0 38.1 45.5 42.3 61.5 52.9
VideoLLaMA3-7B cheng2024videollama 57.6 78.2 29.3 35.7 60.0 59.6 80.8 64.7
InternVL2.5-8B Chen10656429 59.4 83.6 26.7 50.0 60.0 69.2 67.3 68.6
Phi-4-Multimodal-5B abouelenin2025phi 49.7 69.1 26.7 31.0 43.6 57.7 65.4 58.8
Internvideo2.5-8B wang2022internvideo 57.3 85.5 33.3 47.6 56.4 55.8 63.5 64.7
@@
Video-R1-7B videor1 49.2 74.5 22.7 28.6 47.3 48.1 76.9 52.9
@@
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport 54.4 52.9 60.9 35.0 59.5
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport 53.2 52.9 63.0 33.3 51.4
LLaVA-OneVision-7B Li2024LLamavid 48.0 39.7 65.2 35.0 43.2
VideoLLaMA3-7B cheng2024videollama 53.6 45.6 73.9 35.0 50.0
InternVL2.5-8B Chen10656429 60.9 58.8 76.1 43.3 54.1
Phi-4-Multimodal-5B abouelenin2025phi 48.8 55.9 56.5 30.0 40.5
Internvideo2.5-8B wang2022internvideo 59.3 60.3 67.4 43.3 56.8
@@
GPT-4o-mini 60.0 56.0 13.4 58.0 69.4
GPT-4o 70.8 66.7 14.8 65.9 75.5
Gemini-1.5-flash 68.7 64.3 14.8 64.8 77.6
Gemini-2.0-flash 67.0 65.5 14.8 67.0 67.3
@@
Qwen2.5-Omni-7B xu2025qwen25omnitechnicalreport 58.9 53.6 14.1 58.0 63.3
Qwen2.5-VL-7B bai2025qwen25vltechnicalreport 54.0 53.6 12.1 58.0 61.2
LLaVA-OneVision-7B Li2024LLamavid 51.6 33.3 12.1 54.5 34.7
VideoLLaMA3-7B cheng2024videollama 60.3 56.0 13.4 63.6 59.2
InternVL2.5-8B Chen10656429 62.4 57.1 14.8 64.8 69.4
Phi-4-Multimodal-5B abouelenin2025phi 51.9 44.0 9.4 59.1 44.9
Internvideo2.5-8B wang2022internvideo 64.0 60.7 14.1 68.2 67.3
@@ 4.2.1 Overall performance
4.2.2 Closed-source vs open-source models
Our comprehensive evaluation reveals a consistent and significant performance disparity between closed-source and open-source MLLMs across all hierarchical task tiers within the CVBench framework. Quantitative results from Tables 2–4 demonstrate that closed-source models (e.g., GPT-4o, Gemini-2.0-flash) achieve an average accuracy of 67.2% on complex reasoning tasks, substantially outperforming open-source counterparts (e.g., Qwen2.5-VL, LLaVA-OneVision) at 55.8%, representing an 11.4% performance gap. This divergence is particularly pronounced in temporally demanding scenarios such as multi-video temporal reasoning (M.TR), where closed-source models attain 58.7% accuracy compared to merely 22.7% for open-source models, highlighting a critical vulnerability in current publicly available architectures when handling long-range dependencies.
The observed performance chasm stems from fundamental advantages inherent in proprietary systems. Closed-source models leverage superior scalability through larger parameter counts (e.g., GPT-4o’s estimated 1.7 trillion parameters) and more extensive multimodal pretraining datasets, enabling robust context retention across extended video sequences. Furthermore, these models incorporate advanced optimization techniques including specialized attention mechanisms such as cross-video memory buffers and spatiotemporal fusion modules that explicitly model inter-video dependencies, whereas open-source alternatives often rely on adapted single-video frameworks lacking these tailored components. Additionally, commercial systems benefit from substantial computational resources during inference, facilitating sophisticated frame sampling strategies and ensemble methods that remain impractical for most open-source deployments due to hardware constraints.
These findings underscore the necessity for open-source communities to develop specialized architectures specifically designed for cross-video reasoning tasks. The performance gap not only reflects current technological disparities but also highlights the limitations of existing open-source models in integrating information across multiple video sources. Bridging this divide will require innovations in efficient cross-modal attention mechanisms and the development of large-scale multimodal pretraining protocols accessible to broader research communities, ultimately advancing the field toward more equitable and capable multimodal understanding systems.
4.2.3 Sensitivity to input perturbations
Table 5: Ablation on input perturbations.
Run ID #1 #2 #3 Average
@@ 4.2.3 Sensitivity to input perturbations
4.2.4 Frame resolution
We investigate the impact of frame resolution using Qwen2.5-VL, considering different resolutions and varying numbers of input frames, as shown in Table 6. With 8 frames, accuracy increases slightly from 46.0% at 224×224 to 46.8% at 448×448 (+0.8%), indicating that higher resolution leads to better performance. Similar trends are observed with 16 and 32 frames. Additionally, increasing the number of input frames consistently improves performance across both resolutions.
Table 6: Impact of frame number & frame resolution.
@@ 4.2.4 Frame resolution
4.2.5 Number of video instances and duration
Fig. 4 illustrates how video duration and the number of videos affect model performance (Qwen2.5-VL). Accuracy improves with video length up to 900 seconds (48.8% for 0–300s), but declines for longer videos (45.6% for 900–1200s), likely due to increased sequence complexity and context management challenges.
Figure 4: Impact of video number and video duration.
We also highlight the difficulty of multi-video reasoning. Models that perform well on single-video tasks struggle with cross-video tasks. Shown in Fig. 4, the model achieves 52.2%, 47.9%, 44.8% accuracy with two, three, four videos, showing increased complexity of multi-video tasks.
4.2.6 Qualitative analysis of reasoning-based MLLMs

### v4

@@ 1 Introduction
The ability to reason across multiple video streams is fundamental to real-world applications, such as multi-camera surveillance systems, distributed procedural learning, and multi-view activity analysis [Wu2025]. While multimodal large language models (MLLMs) have demonstrated significant progress in single-video understanding [CAI2025111670], their capacity for spatiotemporal pattern reasoning across videos remains a critical and underexplored challenge in pattern recognition research [ZHENG2025111319]. This capability requires models to identify and analyze relational patterns across distinct spatiotemporal contexts [11008449], such as entity correspondences, event sequences, and causal links. The research gap stems not merely from a lack of single-video benchmarks, but from the inherent limitations of existing evaluations in capturing cross-video relational dynamics, which are central to real-world pattern recognition systems.
Current MLLMs face fundamental barriers in cross-video scenarios: 1) Inter-video context linking: models lack mechanisms to persistently integrate and retain object/event states cross-video separated by time and space [Li10658165, 11093290]; 2) Entity disambiguation: overlapping or visually similar entities cross-video lack unique spatiotemporal grounding, causing identification errors; 3) Temporal-causal modeling: existing architectures poorly capture long-range dependencies and causal links spanning asynchronous video sequences [11095110].
@@
Figure 1: Illustration of CVBench task definiton. CVBench focuses on three core tasks that are crucial for real-world applications: (A) cross-video object association (identifying shared entities across different videos), (B) cross-video event association (linking temporal or causal event chains that span across video sequences), and (C) cross-video complex reasoning (integrating commonsense to understand and reason about complex interrelationships between video content). The benchmark format is multi-choice QA, and here we only show the correct answer for visualization.
@@
  • We introduce CVBench, the first diagnostic benchmark explicitly targeting cross-video understanding through object association, event association, and complex reasoning.
  • We release a carefully curated multi-domain dataset with comprehensive annotations, enabling rigorous stress-testing of spatiotemporal integration and multi-view synthesis capabilities.
  • We provide a systematic analysis of state-of-the-art MLLMs on CVBench, exposing critical performance gaps and identifying key research directions for advancing cross-video pattern reasoning in the field.
@@ 2 Related Work
2.1 Multimodal Large Language Models
The evolution of MLLMs has progressed from image-based [Zhang10547418] to video-capable architectures [Fu11093290, ZHANG2026111925], marked by two main improvements: 1) Long-range context modeling that enables efficient handling excessively long temporal contexts [Yue11093104, Song10657734]; 2) Spatio-temporal fusion that integrates both appearance and motion dynamics with language reasoning [Ren10656135, Zhang11094660, Liu11094049]. Despite these advances, current MLLMs remain architecturally constrained to single-video scenario, lacking mechanisms for cross-video interactions, exhibiting inadequate spatiotemporal grounding for overlapping entities, and demonstrating limited modeling of temporal event chains. These fundamental limitations restrict their applicability to multi-camera surveillance and cross-video applications [Yang11037495].
2.2 Video Understanding Benchmarks
A comprehensive and objective benchmark is essential for evaluating MLLMs, enabling the comparison and investigation of the performance of various models [SHAO2026112724]. Recent advances in multimodal learning have led to the development of several benchmarks for video understanding tasks [ZHONG2025111035]. However, existing video evaluation frameworks exhibit critical limitations in assessing cross-video reasoning capabilities. First, general multimodal frameworks are predominantly with broad domain coverage, such as Video-MMMU and Video-MME [Fu11093290]), to evaluate instruction-following capabilities but lacking cross-video relational tasks. Second, long-term video benchmarks contain longer videos with multiple sub-events, however, are confined to single source or similar creators, such as VideoTree [Wang11094563], Video-Bench [Ning11284911], and Chapter-Llama [Ventura11092338]. These deficiencies fundamentally limit their applicability to real-world scenarios in multi-camera surveillance and cross-video procedural learning.
@@ 3.1 Task Definition
  • Object association: Cross-video object recognition, multi-video attribute recognition, joint-video counting, and cross-video entity matching. These require consistent identification of entities across domains or viewpoints despite occlusion or deformation.
  • Event association: Cross-video anomaly detection, scene recognition, key-action recognition, event retrieval, multi-view scene understanding, and temporal reasoning. These demand fusion of spatiotemporal signatures across videos to reconstruct event dynamics.
  • Complex reasoning: Joint-video spatial navigating, video difference captioning , cross-video counterfactual reasoning, joint-video summarization, and procedural transfer. These necessitate higher-order synthesis including spatial imagination, fine-grained reasoning, causal inference and adaptive knowledge transfer.
@@ 3.2 Dataset Collection
Video curation.
We employ task-driven sampling from YouTube to construct diverse video groups (at most four videos per group). Detailed statistic are shown in Fig. 2. We ensure source diversity through the inclusion of multi-creator channels (videos with similar or same creators), cross-domain associations (videos across different domain but with shared topics or intrinsic relations), and the segmentation of long videos (sequential occurrence of event-level activities). Specifically, we measure video quality under the following four curation criteria:
  • Multi-domain coverage, including daily life, sports, surveillance and other commen scenrios;
  • Intrinsic inter-video relationships, such as differences, similarities, co-occurrence, causality causal, temporal order;
  • Duration constraints, where videos are primarily within 10 minutes in length, and those exceeding 10 minutes account for less than 5% of the dataset;
  • Visual quality and availability, requiring videos to have a resolution of at least 720p and a frame rate of 24 fps.
Benchmark #Video #QA #Video Per QA Duration (s) Multi-level CV.U CV.R
TempCompass [tempcompass2024] 410 500 1 11.4
MSVD-QA [xu2017video] 504 13,157 1 9.8
Video-MME [fu2024video] 900 2,700 1 1017.9
NExT-QA [xiao2021next] 1,000 8,564 1 39.5
AutoEval-Video [chen2024autoeval] 327 327 1 14.6
EgoSchema [egoschema2023] 5,063 5,063 1 180.0
ActivityNet-QA [ActivityNet2019] 800 8,000 1 111.4
CVBench (Ours) 1,315 1,000 2 $\sim$ 4 106.6
@@ 3.3 Comparison with Previous Benchmarks
Qwen2.5-Omni-7B [xu2025qwen25omnitechnicalreport] 52.4 70.9 22.7 31.0 56.4 57.7 75.0 60.8
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 51.3 80.0 22.7 26.2 50.9 55.8 69.2 60.8
LLaVA-OneVision-7B [li2024llama] 52.6 83.6 40.0 38.1 45.5 42.3 61.5 52.9
VideoLLaMA3-7B [cheng2024videollama] 57.6 78.2 29.3 35.7 60.0 59.6 80.8 64.7
InternVL2.5-8B [chen2024internvl] 59.4 83.6 26.7 50.0 60.0 69.2 67.3 68.6
Phi-4-Multimodal-5B [abouelenin2025phi] 49.7 69.1 26.7 31.0 43.6 57.7 65.4 58.8
Internvideo2.5-8B [wang2022internvideo] 57.3 85.5 33.3 47.6 56.4 55.8 63.5 64.7
@@
Video-R1-7B [videor1] 49.2 74.5 22.7 28.6 47.3 48.1 76.9 52.9
4 Benchmarked Results
4.1 Experimental Settings
Models.
Our evaluation comprehensively assesses various MLLMs, including closed-source and open-source models. A prominent model in our suite is Gemini [team2023gemini], a closed-source model known for its sophisticated multimodal processing capabilities. On the open-source side, Qwen2.5-VL [videor1] stands out, offering transparency with its accessible codebase for further exploration and customization. We also evaluate advanced multimodal reasoning models such as Video-R1 [videor1], which leverages reinforcement learning to integrate multimodal information and perform complex cross-video reasoning tasks. For ablation study, we choose Qwen2.5VL-7B as the default model.
Settings.
@@ 4.1 Experimental Settings
Qwen2.5-Omni-7B [xu2025qwen25omnitechnicalreport] 54.4 52.9 60.9 35.0 59.5
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 53.2 52.9 63.0 33.3 51.4
LLaVA-OneVision-7B [li2024llama] 48.0 39.7 65.2 35.0 43.2
VideoLLaMA3-7B [cheng2024videollama] 53.6 45.6 73.9 35.0 50.0
InternVL2.5-8B [chen2024internvl] 60.9 58.8 76.1 43.3 54.1
Phi-4-Multimodal-5B [abouelenin2025phi] 48.8 55.9 56.5 30.0 40.5
Internvideo2.5-8B [wang2022internvideo] 59.3 60.3 67.4 43.3 56.8
@@
GPT-4o-mini 60.0 56.0 60.4 58.0 69.4
GPT-4o 70.8 66.7 74.5 65.9 75.5
Gemini-1.5-flash 68.7 64.3 70.5 64.8 77.6
Gemini-2.0-flash 67.0 65.5 67.8 67.0 67.3
@@
Qwen2.5-Omni-7B [xu2025qwen25omnitechnicalreport] 58.9 53.6 61.1 58.0 63.3
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 54.0 53.6 49.7 58.0 61.2
LLaVA-OneVision-7B [li2024llama] 51.6 33.3 65.8 54.5 34.7
VideoLLaMA3-7B [cheng2024videollama] 60.3 56.0 61.1 63.6 59.2
InternVL2.5-8B [chen2024internvl] 62.4 57.1 61.7 64.8 69.4
Phi-4-Multimodal-5B [abouelenin2025phi] 51.9 44.0 54.4 59.1 44.9
Internvideo2.5-8B [wang2022internvideo] 64.0 60.7 62.4 68.2 67.3
@@ 4.2 Results and Findings
Closed-source vs. open-source models.
Closed-source models consistently outperform open-source models. This suggests that proprietary architectures may utilize more refined optimization techniques or superior capabilities in handling complex cross-video relationships, revealing a substantial gap between commercial-grade and publicly available MLLMs in this domain.
Run ID #1 #2 #3 Avg.
@@
Table 5: Ablation on input perturbations.
Sensitivity to input perturbations.
@@
Table 6: Impact of frame number & resolution.
Frame resolution.
We investigate the impact of frame resolution using Qwen2.5-VL, considering different resolutions and varying numbers of input frames, as shown in Table 6,. With 8 frames, accuracy increases slightly from 46.0% at 224×224 to 46.8% at 448×448 (+0.8%), indicating that higher resolution leads to better performance. Similar trends are observed with 16 and 32 frames. Additionally, increasing the number of input frames consistently improves performance across both resolutions.
Figure 4: Impact of video number and duration.
Number of video instances & duration.
Fig. 4 illustrates how video duration and the number of videos affect model performance (Qwen2.5-VL). Accuracy improves with video length up to 900 seconds (48.8% for 0–300s), but declines for longer videos (45.6% for 900–1200s), likely due to increased sequence complexity and context management challenges. We also highlight the difficulty of multi-video reasoning. Models that perform well on single-video tasks struggle with cross-video tasks. Shown in Fig. 4, the model achieves 52.2%/47.9%/44.8% accuracy with two/three/four videos, showing increased complexity of multi-video tasks.
Qualitative analysis of reasoning-based MLLMs.